OpenStudy (anonymous):

Which shows this equation in standard form? 8 – 4(x – y) = 2x + 6 A. 6x – 4y = 2 B. 6x – 4y = –2 C. 4x – 6y = 2 D. 6x + 4y = 2

OpenStudy (anonymous):

8-4x+4y=2x+6 -4x-2x+4y=6-8 -6x+4y=-2 6x-4y=2 <---B

OpenStudy (anonymous):

I'm going to say D because, look, it's the only different problem. it's Ax+By=c

OpenStudy (anonymous):

@whalexnuker No.. try out making ur problem !@! Answer will be B

OpenStudy (anonymous):

ty and @whalexnuker dont nuke whales

OpenStudy (anonymous):

lol i dont care if u do

OpenStudy (anonymous):

Hahahaaa... :P